Patel Foundry in Youngstown, Ohio, uses a predetemmined manufacturing E(Clck the icon to view the costs) overhead rate to allocate ovehead to individual jobs based on the machine hours requred Al the beginning of the year, the company expected to incar the tollowing 1. Compute Patets predetermined manufacturing overhead rate 2. How much manufacturing overhead was allocated to jobs during the year? . How much manufactuing overhead was incunmed during the year?is manufacturing overhead underallocated or overallocated at the end of the year? By how much? 4. Were the jobs overcosted or undercosted? By how much Requirement 1. Compute Pates predetermined manufacturing overhead rate Determine the formula to calculiate the predetermined overhead rate then caladate the rate Estimated yearly direct iabor cot Estmated yearly dred labor hour Estmated yearly machine hours Estmated yearly overhead costs
Data Table Manufacturing overhead costs Direct labor cost Machine hours. S 600,000 S 1,650,000 75,000 At the end of the year, the company had actually incurred the following: Direct labor cost Depreciation on manufacturing plant and $ 1,170,000 495,000 18,500 24,500 16,500 9000 equipment Property taxes on plant Sales salaries Delivery drivers wages Plant janitors wages Machine hours 53,500 hours Print Done
Answer #1

1) Predetermine overhead rate =

Estimated manufacturing overhead / Estimated machine hour = Predetermine overhead rate
600000 / 75000 = $8 per machine hour

2) Manufacturing overhead allocated = 53500*8 = $428000

3) Actual manufacturing overhead = 495000+18500+9000 = $522500

Under applied overhead by (522500-428000) = $94500

4) Job is under costed by $94500